OpenStudy (anonymous):

The volume of a cube is 512 cm3. What would change if the volume of the cube were measured in cubic meters?

OpenStudy (jdoe0001):

 so what's "x" well \(\bf volume = 512\qquad volume=x^3\\ \quad \\ 512=x^3\implies \sqrt[3]{512}=x\) what's "x"?